OpenStudy (anonymous):

if 1 equivalent = 10000 milliequivalent (abbr : meq). If 53g Na2CO3 put into a volumetric flask of 250ml, dissolved by adding water until the linemark. the normality of this solution is?

OpenStudy (abhisar):

okay lets first calculate equivalent mass of Na2CO3

OpenStudy (abhisar):

I think u knw that Na2CO3 is a salt ?

OpenStudy (anonymous):

yaa i know

OpenStudy (abhisar):

okay that's great ! Now equivalent weighgt for salt = Molecular weight/Total charge on cation/anion

OpenStudy (abhisar):

Here total charge on cation or anion is 2

OpenStudy (abhisar):

and we just calculated the molecular weight i.e 106

OpenStudy (abhisar):

so eq wt = 53

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

ya i understand!!

OpenStudy (abhisar):

now the next step is to calculate the equivalents of Na2CO3 in 53g Na2CO3

OpenStudy (abhisar):

It will be mass of solute/ equivalent mass=53/53=1

OpenStudy (abhisar):

Final step will be Normality = equivalents/volume(L)= (1X1000)/250 = 4

OpenStudy (anonymous):

perfect thank u soo much!! i really understand to the bit :)

OpenStudy (abhisar):

\(\Huge\text{Anytime !}\) \(\huge\ddot\smile\)